How many milliliters of 3 F ammonia can be prepared from 100ml of concentrated ammonia (28.4 per cent NH3 by weight, having a density of 0.808 grams per ml)?


1
Expert's answer
2021-03-04T06:08:03-0500

1) Molar concentration of the given ammonia solution is:

CM=n/V

n=m/M

M (NH3) = 18 g/mol

Lets assume that we have 1 L of NH3 solution.

density = m/V

m = density x V

m (NH3)solution = 0.808 x 1 = 0.808 kg

%= (100 x m(solute))/(m(solute)+m(solvent))

m(NH3) = (28.4 x 808)/100 = 229.5 g

n(NH3) = 229.5/18 = 12.8 mol

CM(NH3) = 12.8/1 = 12.8 mol/L


2) 100 ml of 12.8 molar concentration is equal to: 0.1 x 12.8 = 1.28 moles

V (NH3) = n/CM = 1.28/3 = 0.43 L